OpenStudy (anonymous):

tan(theta)+cot(theta)=csc(theta)sec(theta) i know this is a proof but can you show how to use identities to make the sides look the same?

OpenStudy (anonymous):

Have you tried converting everything to sin and cos?

OpenStudy (jdoe0001):

\(\bf tan(\theta)+cot(\theta)\implies \cfrac{sin(\theta)}{cos(\theta)}+\cfrac{cos(\theta)}{sin(\theta)}\implies ?\)